# A Comprehensive Survey on Uncertainty Quantification for Deep Learning

### v4

@@ 1. Introduction
Deep neural network (DNN) models have achieved remarkable success in computer vision, natural language processing, as well as science and engineering domains (Pandey et al., 2022; LeCun et al., 2015; Deng et al., 2009). Existing popular DNN models can be largely regarded as a deterministic function, which maps the input features to a target prediction through hierarchical representation learning (Bengio et al., 2013). While these DNN models often achieve strong performance in overall prediction accuracy, it is also well-recognized that they sometimes make unexpected, incorrect, but overconfident predictions, especially in a complex real-world environment (Reichstein et al., 2019). This can cause serious consequences in high-stake applications, such as autonomous driving (Choi et al., 2019), medical diagnosis (Begoli et al., 2019), and disaster response (Alam et al., 2017). In this regard, a DNN model should be aware of what it does not know in order to avoid overconfident predictions. For example, in the medical domain, when DNN-based automatic diagnosis systems encounter uncertain cases, a patient should be referred to a medical expert for more in-depth analysis to avoid fatal mistakes. In an autonomous vehicle, if a DNN model knows in what scenarios (e.g., bad weather) it tends to make mistakes in estimating road conditions, it can warn the driver to take over and avoid potential crashes.
”Knowing what a DNN model does not know” comes down to placing appropriate uncertainty scores in its predictions, also called uncertainty quantification (UQ). Uncertainty in DNNs may come from different types of sources, including data uncertainty and model uncertainty (Yarin, 2016). Data uncertainty (also aleatory uncertainty) is an inherent property of the data, which originates from the randomness and stochasticity of the data (e.g., sensor noises) or conflicting evidence between the training labels (e.g., class overlap). Data uncertainty is often considered irreducible because we cannot reduce it by adding more training samples. Model uncertainty (also epistemic uncertainty), on the other hand, comes from the lack of evidence or knowledge during model training or prediction for a new test sample, e.g., limited training samples, sub-optimal DNN model architecture or parameter learning, and out-of-distribution (OOD) samples.
@@
2. Problem definition
Given training data $\mathcal{D}_{tr}=\{\boldsymbol{x}_{i},{y}_{i}\}_{i=1}^{n}$ with input features $\mathbf{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{n}$, $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ and the target labels $\mathbf{Y}=\{{y}_{i}\}_{i=1}^{n}$, the learning problem is to find a parameterized neural network model $y=f(\boldsymbol{x};\boldsymbol{\boldsymbol{\theta}})$ that fits we on the training data (and potentially generalizes well on the unseen test data). The model is learned by minimizing the empirical risk, defined as the average loss over the training data: (Murphy, 2012):
(1) $\operatorname*{arg\,min}_{f}R(\mathcal{D}_{tr},f),\ \text{where}\ R(\mathcal{D}_{tr},f)=\frac{1}{n}\sum_{i=1}^{n}{l}(y_{i},f(\boldsymbol{x}_{i}))$
In addition to making an accurate prediction, the goal of UQ is to also obtain a parametric uncertainty framework $g$ that outputs an uncertainty estimate $u=g(\boldsymbol{x};\boldsymbol{\phi})\in\mathbb{R^{+}}$ reflecting the confidence of the model prediction. To this end, we formulate the problem as an optimization problem that minimizes a new loss function, which is a combination of the empirical risk and the expected calibration error (ECE) (Nixon et al., 2019):
(2) $\small\begin{split}\operatorname*{arg\,min}_{f,g}\hat{R}(\mathcal{D},f,g),\ \text{where}\ \hat{R}(\mathcal{D},f,g)=R(\mathcal{D}_{tr},f)+\text{ECE}(\mathcal{D}_{tr},f,g)\end{split}$
ECE measures the consistency between the prediction error and the uncertainty of the prediction. Specifically, the uncertainty interval is grouped into fixed bins, and the average of the difference between the uncertainty and error in each bin is compared. ECE encourages higher uncertainty on larger error predictions.
Given a sample $\boldsymbol{x}$, the goal is to predict both the target and associated uncertainty $(y,u)=(f_{\boldsymbol{\theta}}(\boldsymbol{x}),g_{\boldsymbol{\phi}}(\boldsymbol{x}))$. Prediction and uncertainty estimation are listed as two separate modules, $f$ and $g$. However, some approaches may represent both in a single, coherent framework.
@@ 3. Types of uncertainty source
This section provides the background of two major types of uncertainty: data uncertainty and model uncertainty. We discuss the potential sources of each type and its representation.
@@ 3.1.1. Source of data uncertainty
Data uncertainty (aka. aleatoric uncertainty) arises from the data’s inherent randomness, noise, or overlapping feature distribution in different classes, which is irreducible even with more training data (Yarin, 2016). Randomness or noise in data can occur in data acquisition. The uncertainty in the data acquisition process is due to instrument errors, inappropriate frequency of data acquisition, and data transmission errors (Hariri et al., 2019). For example, certain data acquisition devices cannot work properly in a complex environment (e.g., poor weather conditions). Factors such as inappropriate sampling methods, data storage, data representation methods, and interpolation techniques also affect uncertainty during data processing  (Yarin, 2016).
For example, for spatiotemporal data, collected from various space and airborne platforms (e.g., CubeSat, UAVs), the data uncertainty may result not only from the sensor errors associated with the data acquisition devices but also from the fact that data is acquired in a digital format (which is discrete in nature) (Cheng et al., 2014) even though the underlying phenomenon is continuous. The uncertainty of the representation of an object’s movement is also affected by the frequency with which location samples are taken, i.e., the sample rate (Pfoser and Jensen, 1999). Uncertainty in the data can also accumulate from multiple sources and may be propagated into the model.
@@ 3.1.2. Data uncertainty representation
Consider a training dataset $\mathcal{D}_{tr}$ drawn from the distribution ${p_{tr}(\boldsymbol{x},y)}$ over the $d$ dimensional input features $\boldsymbol{x}\in\mathbb{R}^{d}$ and target class labels $y\in\{\omega_{1},...,\omega_{k}\}$ for classification problem with $k$ classes or $y\in\mathbb{R}^{k}$ for regression model. In the context of a discriminative classification task, the uncertainty of the class variable $y$ given a specific input instance $\boldsymbol{x}$ is defined as the entropy of the true condition class distribution $\mathcal{H}[p_{tr}(y|\boldsymbol{x})]$ as Eq. 3 shows. The conditional entropy describes the randomness of the class distribution due to the overlap of feature values among samples in different classes.
(3) $\mathcal{H}[p_{tr}(y|\boldsymbol{x})]=-\sum_{i=1}^{k}p_{tr}(y=\omega_{i}|\boldsymbol{x})\log(p_{tr}(y=\omega_{i}|\boldsymbol{x}))$
@@
For classification problems, data uncertainty arises from the natural complexity of the data and the structure of the class boundary on the feature space. Take a toy distribution in Fig. 2 as an example, which consists of two normally distributed clusters. Each cluster (color) represents a separate class. The dataset in Fig. 2 (a) has clearer class boundary (lower data uncertainty). The entropy of most samples is low except for those near the class boundary, as in Fig. 2 (b). In contrast, the dataset in Fig. 2 (c) has more overlap between two classes, corresponding to higher data uncertainty in Fig. 2 (d).
@@
Unlike classification problems, for the regression problem data uncertainty arises from the inherent noise (variability) in the data generation or collection process:$y=f(\boldsymbol{x})+\epsilon(\boldsymbol{x})$, where $\epsilon(\boldsymbol{x})$ is the observation or measurement noise in the dataset. There are two classes of noise: homeostatic noise and heteroscedastic noise (Kendall and Gal, 2017). Homeostatic noise assumes a constant observation noise over all the inputs $\boldsymbol{x}$. Heteroscedastic noise, on the other hand, models the observation noise as a function of input $\epsilon(\boldsymbol{x})\sim p(\epsilon|\boldsymbol{x})$ (e.g., heteroscedastic Gaussian noise). The heteroscedastic noise model is useful in the case where the noise level vary for different samples.
3.2. Model uncertainty
3.2.1. Source of model uncertainty
The model uncertainty (aka Epistemic uncertainty) represents the uncertainty in the models’ predictions related to the imperfect model training process (e.g., due to a lack of knowledge on model choices based on the current input data). It is reducible given more training data. Model uncertainty can be divided into three types: uncertainty in model architecture; uncertainty in the model parameters; uncertainty due to a mismatch between training and test dataset distribution. For the first type, model architecture uncertainty arises due to the lack of understanding of which type of model architecture is most suitable for the data. For example, for deep learning models, there exists uncertainty regarding how many neural network layers, and how many neurons in each layer are optimum for the given dataset. A too-complex model can cause overfitting. Uncertainty in model parameters is due to the unknown optimal parameters and it may arise from an improper training strategy, lack of training instances, or local optimal solution. There is no guarantee that the converged weight values correspond to the global minimum of the loss function. The last type of model uncertainty is caused by data distribution drift, which means the test sample distribution is different from the training distribution. The issue is not uncommon for real-world deployment of deep learning models, also known as out-of-distribution(OOD) data (Schwaiger et al., 2020), as real-world test cases can be very complex.
3.2.2. Model uncertainty representation
Compared with data uncertainty, model uncertainty is more difficult to estimate since the source of model uncertainty can arise from three different aspects: model parameters, model architectures, and dataset distribution shift. In general, different methods are adopted to represent the model uncertainty from each type. First, uncertainty from model parameters can be estimated with a Bayesian neural network (BNN) (Jospin et al., 2022). BNN assumes a prior over the model parameters and aims to infer the posterior distribution of model parameters to reflect the parameters uncertainty. Second, uncertainty arising from the model architectures are estimated with ensembles of neural network (deep ensembles). The intuition is to construct an ensemble of neural network architectures and each model is trained separately. The predictions of the ensemble on an input form a distribution on the target variable. Thus the variance of the target variable predictions can be an estimation of the prediction uncertainty. The third type is the uncertainty from dataset distribution shift, which is caused by the mismatch between the training dataset distribution and that of the test samples. The further away a new test sample is from the training samples, the greater model uncertainty there is.
@@ 3.2.2. Model uncertainty representation
In summary, we highlight two primary types of uncertainty and discuss their potential sources and representations. As Fig. 3 shows, data uncertainty arises from its inherent property of the input data features and labels, and model uncertainty stems from the misspecification of model architectures, parameters, and the dataset distribution shit. Depending on the nature of the application, the predominant type of uncertainty may vary, and specific methods may be necessary to address them. These uncertainties underscore the importance of robust analysis and interpretation of data and model outputs in various fields.
Figure 3. Two types of uncertainty source
@@ 4. Application domains
Medical diagnosis: Deep neural network models have achieved tremendous success in a wide variety of medical applications, including medical imaging, clinical diagnosis support, and treatment planning (Oh et al., 2020; Van Tulder and De Bruijne, 2016; Cireşan et al., 2013). However, one critical concern is that deep learning models tend to be over-confident even for a wrong prediction (Loftus et al., 2022), causing serious consequences. Thus, we need to estimate the prediction uncertainty (confidence). Both data uncertainty and model uncertainty exist in medical problems. Data uncertainty comes from noisy measurements from medical devices, ambiguous class labels (e.g., non-consensus tumor boundary annotations on imagery between different radiologists), and registration errors between medical imagery taken at different times or from different devices (Gong et al., 2022). Model uncertainty also exists because patient subjects in the test cases may not be well-represented in the training set. There exist several challenges to developing UQ methods. First, there is diverse noise or uncertainty in medical data. Second, interpretability in uncertainty quantification is important, which is not well solved in medical problems. Existing uncertainty-aware deep learning models in medical domains can be categorized into those related to medical imaging and those for non-medical imaging applications (Loftus et al., 2022). In medical imaging, deep learning is often used for segmentation or classification for MRI, ultrasound, and cohere tomography (CT) imagery (Kohl et al., 2018; Edupuganti et al., 2020). These studies often focus on data uncertainty due to ambiguous labels (Natekar et al., 2020; Qin et al., 2021), or image registration uncertainty (Chen et al., 2021a). Non-medical imaging applications are mostly related to clinical diagnosis support and treatment planning from Electronic Health Records (EHR). Recurrent neural networks and graph neural networks have been applied to EHR data to predict patient diseases and possible treatment (Zhang et al., 2020). The presence of significant variability in personalized predictions (Dusenberry et al., 2020) requires a model to capture prediction uncertainty.
@@
Geoscience: With the advance of GPS and remote sensing technologies, a growing volume of spatiotemporal data are collected from spaceborne, airborne, seaborne, and terrestrial platforms (Shekhar et al., 2015; NASA Goddard, 2019; Florida and Wildlife, 2021). The emerging spatiotemporal big data, increasing computational power (GPUs), and the recent advances in deep learning technologies provide unique opportunities for advancing our knowledge about the Earth system (Reichstein et al., 2019). For example, deep learning has been used to predict river flow and temperature (Jia et al., 2021) and hurricane tracks (Kim et al., 2019). Uncertainty quantification for deep learning is important in geoscience because of the high-stake decision-making involved (e.g., evacuation planning based on hurricane tracking with a “cone of uncertainty”). Several challenges exist due to the unique characteristics of spatiotemporal data. First, spatiotemporal data have various spatial, temporal, and spectral resolutions and diverse sources of noise and errors (e.g., sensor noise, obstacles, and atmospheric effects in remote sensing signals (Licata and Mehta, 2022), GPS errors). Second, spatial registration error and uncertainty may exist when co-registering different layers of geospatial data into the same spatial reference system (He et al., 2022). Third, spatiotemporal data are heterogeneous, i.e., the data distribution often varies across different regions or time periods (Jiang et al., 2019). Thus, a deep learning model trained in one region (or time) may not generalize well to another region (or time). This issue is particularly important when spatial observation samples are sparsely distributed, causing uncertainty when inferring the observations at other locations in continuous space (Hengl et al., 2017).
@@
Transportation: Deep learning for transportation data (e.g., ground sensors and video cameras on the road) provide unique opportunities to monitor traffic conditions, analyze traffic patterns and improve decision-making. For instance, temporal graph neural networks are used to predict traffic flows(e.g., congestion or accidents) (Zhao et al., 2019; Mo and Fu, 2022). Deep learning plays a critical role in autonomous driving (e.g, using lidar sensors and optical cameras to detect road lanes and other vehicles or pedestrians). Uncertainty quantification for AI in transportation is challenging due to the temporal dynamics, complex road environment, and the existence of noise and uncertainty (e.g., omission, sparse sensor coverage, errors, or inherent bias). For example, highly crowded events can disrupt normal traffic flows on road networks. Existing works on trajectory prediction uncertainty consider the data uncertainty due to the sparse or insufficient training data (Zhou et al., 2022), and erroneous or missing measurements due to signal loss (Markos et al., 2021). Other works consider complex environment factors such as extreme weather (Pang et al., 2021; Zhu et al., 2022; Lempert et al., 2022; Wang et al., 2019a). Uncertainty in short-term traffic status forecasting (such as volume, travel speeds, and occupancy) is related to the stochastic environment and model training (Wang et al., 2014), while uncertainty in long-term traffic modeling comes from the exogenous factors of traffic flow (e.g., rainstorms and snowstorms) (Li et al., 2022).

### v5

# A Survey on Uncertainty Quantification Methods for Deep Learning

@@ 1. Introduction
Deep neural network (DNN) models have achieved remarkable success in computer vision, natural language processing, as well as science and engineering domains (Pandey et al., 2022; LeCun et al., 2015; Deng et al., 2009). Existing popular DNN models can be largely regarded as a deterministic function, which maps the input features to a target prediction through hierarchical representation learning (Bengio et al., 2013). While these DNN models often achieve strong performance in overall prediction accuracy, it is also well-recognized that they sometimes make unexpected, incorrect, but overconfident predictions, especially in a complex real-world environment (Reichstein et al., 2019). This can cause serious consequences in high-stake applications, such as autonomous driving (Choi et al., 2019), medical diagnosis (Begoli et al., 2019), and disaster response (Alam et al., 2017).
In this regard, a DNN model should be aware of what it does not know in order to avoid overconfident predictions. For example, in the medical domain, when DNN-based automatic diagnosis systems encounter uncertain cases, a patient should be referred to a medical expert for more in-depth analysis to avoid fatal mistakes. In an autonomous vehicle, if a DNN model knows in what scenarios (e.g., bad weather) it tends to make mistakes in estimating road conditions, it can warn the driver to take over and avoid potential crashes.
"Knowing what a DNN model does not know" comes down to placing appropriate uncertainty scores in its predictions, also called uncertainty quantification (UQ). Uncertainty in DNNs may come from different types of sources, including data uncertainty and model uncertainty (Yarin, 2016). Data uncertainty (also aleatory uncertainty) is an inherent property of the data, which originates from the randomness and stochasticity of the data (e.g., sensor noises) or conflicting evidence between the training labels (e.g., class overlap). Data uncertainty is often considered irreducible because we cannot reduce it by adding more training samples. Model uncertainty (also epistemic uncertainty), on the other hand, comes from the lack of evidence or knowledge during model training or prediction for a new test sample, e.g., limited training samples, sub-optimal DNN model architecture or parameter learning, and out-of-distribution (OOD) samples.
@@
2. Supervised Learning Problem definition
Given training data $\mathcal{D}_{tr}=\{\boldsymbol{x}_{i},{y}_{i}\}_{i=1}^{n}\subset\mathcal{X}\times\mathcal{Y}$, $\mathcal{X}$ is the input sample space, and $\mathcal{Y}$ is the target variable space for the input sample. Each training instance is assumed to be independent and identically distributed (i.i.d.) from some unknown probability distribution $p(\boldsymbol{x},y)$ on the space $\mathcal{X}\times\mathcal{Y}$. Given a hypothesis space $\mathcal{H}$ consisting of hypotheses $h:\mathcal{X}\rightarrow\mathcal{Y}$ and a loss function $l$ that measures the discrepancy between prediction and ground-truth, a learning problem aims to find the best hypothesis in the hypothesis space that minimizes the loss (Hüllermeier and Waegeman, 2021; Bengs et al., 2023):
[…]
In practice, the model is fitted by minimizing the empirical risk (Murphy, 2012), defined as the average loss over the training data $\mathcal{D}_{tr}$:
(2) $\tilde{h}=\operatorname*{arg\,min}_{h\in\mathcal{H}}R_{emp}(\mathcal{D}_{tr},h),\ \text{where}\ R_{emp}(\mathcal{D}_{tr},h)=\frac{1}{n}\sum_{i=1}^{n}{l}(y_{i},h(\boldsymbol{x}_{i}))$
@@ 3. Types of uncertainty source
Based on the mathematical formulation of supervised learning in Section 2, we provide the definitions for two major types of uncertainty, data uncertainty and model uncertainty. We discuss the sources of each type and their representations.
3.1. Model uncertainty
3.1.1. Source of model uncertainty
The model uncertainty (aka epistemic uncertainty) represents
Figure 2. Visualization on various model uncertainty sources.
[…]
Table 3.1.1 summarizes different sources of model uncertainty in the supervised learning framework. The first type, model architecture uncertainty is due to the lack of knowledge of which type of model architecture is most suitable for the data. Because of this, the optimal solution $h^{*}$ selected from $\mathcal{H}$
Table 1. The corresponding supervised learning notation for different model uncertainty sources
Model uncertainty
source
Corresponding notation
in supervised learning
framework
Model
architecture
Optimal solution $h^{*}$ in
$\mathcal{H}$ does not align
with theoretical
optimal $f^{*}$ in $\mathcal{F}$
Model
parameters
Learned model $\tilde{h}$
does not align with
optimal $h*$ in $\mathcal{H}$
Sample
distribution
Theoretical optimum
$f_{1}^{*}$ and $f_{2}^{*}$ mismatch
under different sample
distribution $p(\boldsymbol{x},y)$
under specific architecture (hypothesis space) is different from the theoretical optimal $f^{*}$. This is true even if the learning algorithm is perfect. Second, due to limited training data or imperfect learning algorithms in the optimization process, the learned optimal solution $\hat{h}$ with the empirical training loss may not align to the optimal solution $h^{*}$ with the specific architecture $\mathcal{H}$. This leads to model uncertainty related to model parameter learning. Third, model uncertainty can be related to the (training) sample distribution. For instance, learning samples may come from different distributions, e.g., $p_{1}(\boldsymbol{x},y)$ and $p_{2}(\boldsymbol{x},y)$. Thus, the theoretical optimal solution $f^{*}$ varies for different population distributions, causing uncertainty in model predictions on out-of-distribution samples. In addition, when a model makes predictions on a test sample that is far away from other training samples (or surrounded by sparse training samples) in the feature space $\mathcal{X}$ (see Fig. 7), the prediction tends to have higher uncertainty. Note that the three types of model uncertainty sometimes overlap with each other. Model uncertainty related to sample distribution is also connected to uncertainty in parameter learning. For example, model uncertainty due to the lack of training samples near a test sample can also be considered as model parameter learning uncertainty.
3.1.2. Model uncertainty representation
As discussed above, the source of model uncertainty can arise from three different aspects: model parameters, model architectures, and the sample distribution. In general, different methods are adopted to represent the model uncertainty from each type. First, uncertainty from model parameters is from the lack of knowledge of optimal parameters for the given problem. It can be estimated with a Bayesian neural network (BNN) (Jospin et al., 2022). BNN assumes a prior over the model parameters and aims to infer the posterior distribution of model parameters to reflect the parameter uncertainty. Second, uncertainty arising from the model architectures is due to the lack of knowledge regarding the optimal hypothesis space. It can be estimated with ensembles of neural networks (deep ensembles). The intuition is to construct an ensemble of neural network architectures and each model is trained separately. The predictions of the ensemble on an input form a distribution on the target variable. Thus the variance of the target variable predictions can be an estimation of the prediction uncertainty. The third type is the uncertainty from sample distribution, which is caused by the mismatch between the training dataset distribution and that of the test samples. More details of the categorization and methods are discussed in Section 5.
@@ 3.2.1. Source of data uncertainty
Data uncertainty (aka aleatoric uncertainty) arises from the data’s inherent randomness, noise, or overlapping feature distribution in different classes, which is irreducible even with more training data (Yarin, 2016). Randomness or noise in data can arise in data acquisition. The uncertainty in the data acquisition process is due to instrument errors, inappropriate frequency of data acquisition, and data transmission errors (Hariri et al., 2019). For example, certain data acquisition devices cannot work properly in a complex environment (e.g., poor weather conditions). Inappropriate sampling methods, data storage, data representation methods, and interpolation techniques also affect uncertainty during data processing (Yarin, 2016).
For example, for spatiotemporal data collected from various space and airborne platforms (e.g., CubeSat, UAVs), the data uncertainty may result not only from the sensor errors associated with the data acquisition devices but also from the fact that data are acquired in a digital format (which is discrete in nature) (Cheng et al., 2014) even though the underlying phenomenon is continuous. The uncertainty of the representation of an object’s movement is also affected by the frequency with which location samples are taken, i.e., the sample rate (Pfoser and Jensen, 1999). Uncertainty in the data can also accumulate from multiple sources and may be propagated into the model.
@@ 3.2.2. Data uncertainty representation
Consider a training dataset $\mathcal{D}_{tr}$ drawn from the distribution ${p(\boldsymbol{x},y)}$ over the $d$ dimension input features $\boldsymbol{x}\in\mathbb{R}^{d}$ and target class labels $y\in\{\omega_{1},...,\omega_{k}\}$ for a classification problem with $k$ classes or $y\in\mathbb{R}$ for a regression model. There are several techniques for data uncertainty representation to account for the inherent randomness between the mapping from $\boldsymbol{x}$ to $y$. In the context of a discriminative classification task, one method to represent the uncertainty of the class variable given a specific input $\boldsymbol{x}$ is the maximum class probability $\max p(y|\boldsymbol{x})$. Another method uses the entropy of the condition class distribution $p(y|\boldsymbol{x})$, which describes the randomness of the class distribution due to the overlap of feature values among samples in different classes. One other method is set-based methods, e.g., conformal prediction (Shafer and Vovk, 2008), which aims to ensure the correct answer in the prediction set with a predefined confidence level.
@@
Data uncertainty arises from the data’s natural complexity and the class boundary structure on the feature space. Take a toy distribution in Fig. 3 as an example, which consists of two normally distributed clusters. Each cluster (color) represents a separate class. The dataset in Fig. 3 (a) has a clearer class boundary (lower data uncertainty). The entropy of most samples is low except for those near the class boundary, as in Fig. 3 (b). In contrast, the dataset in Fig. 3 (c) has more overlap between the two classes, corresponding to higher data uncertainty in Fig. 3 (d).
@@
Beyond the inherent complex class boundary, data uncertainty can also come from the inherent noise (variability) in the data generation or collection process. For example, for a regression problem, the observations can be represented by: $y=f(\boldsymbol{x})+\epsilon(\boldsymbol{x})$, where $f(\boldsymbol{x})$ is the true data generation function, and $\epsilon(\boldsymbol{x})$ represents the measurement noise in the dataset. There are two classes of noise: homoscedastic and heteroscedastic noise (Kendall and Gal, 2017). Homoscedastic noise assumes a constant observation noise over all the inputs $\boldsymbol{x}$. Heteroscedastic noise, on the other hand, models the observation noise as a function of input $\epsilon(\boldsymbol{x})\sim p(\epsilon|\boldsymbol{x})$ (e.g., heteroscedastic Gaussian noise). The heteroscedastic noise model is useful in the case where the noise level varies for different samples.
3.3. Summary
@@
In summary, we highlight two primary types of uncertainty and discuss their potential sources and representations. As Fig. 4 shows, data uncertainty arises from the inherent property of the input data features and labels, and model uncertainty stems from the misspecification of model architectures, parameters, and the dataset distribution shift. Depending on the nature of the application, the predominant type of uncertainty may vary, and specific methods may be necessary to address them. These uncertainties underscore the importance of robust analysis and interpretation of data and model outputs in various fields.
Figure 4. Different types of uncertainty source
@@ 4. Application domains
Medical diagnosis: Deep neural network models have achieved tremendous success in a wide variety of medical applications, including medical imaging, clinical diagnosis support, and treatment planning (Oh et al., 2020; Van Tulder and De Bruijne, 2016; Cireşan et al., 2013; Cao et al., 2024). However, one critical concern is that deep learning models tend to be over-confident even for a wrong prediction (Loftus et al., 2022), causing serious consequences. Thus, we need to estimate the prediction uncertainty (confidence). Both data uncertainty and model uncertainty exist in medical problems. Data uncertainty comes from noisy measurements from medical devices, ambiguous class labels (e.g., non-consensus tumor boundary annotations on imagery between different radiologists), and registration errors between medical imagery taken at different times or from different devices (Gong et al., 2022). Model uncertainty also exists because patient subjects in the test cases may not be well-represented in the training set. There exist several challenges to developing UQ methods. First, there is diverse noise or uncertainty in medical data. Second, interpretability in uncertainty quantification is important, which is not well solved in medical problems. Existing uncertainty-aware deep learning models in medical domains can be categorized into those related to medical imaging and those for non-medical imaging applications (Loftus et al., 2022). In medical imaging, deep learning is often used for segmentation or classification for MRI, ultrasound, and coherence tomography (CT) imagery (Kohl et al., 2018; Edupuganti et al., 2020). These studies often focus on data uncertainty due to ambiguous labels (Natekar et al., 2020; Qin et al., 2021), or image registration uncertainty (Chen et al., 2021a). Non-medical imaging applications are mostly related to clinical diagnosis support and treatment planning from Electronic Health Records (EHR). Recurrent neural networks and graph neural networks have been applied to EHR data to predict patient diseases and possible treatment (Zhang et al., 2020). The presence of significant variability in personalized predictions (Dusenberry et al., 2020) requires a model to capture prediction uncertainty.
@@
Geoscience: With the advance of GPS and remote sensing technologies, a growing volume of spatiotemporal data is collected from spaceborne, airborne, seaborne, and terrestrial platforms (Shekhar et al., 2015; NASA Goddard, 2019; Florida and Wildlife, 2021). The emerging spatiotemporal big data, increasing computational power (GPUs), and the recent advances in deep learning technologies provide unique opportunities for advancing our knowledge about the Earth system (Reichstein et al., 2019). For example, deep learning has been used to predict river flow and temperature (Jia et al., 2021) and hurricane tracks (Kim et al., 2019). Uncertainty quantification for deep learning is important in geoscience because of the high-stakes decision-making involved (e.g., evacuation planning based on hurricane tracking with a “cone of uncertainty”). Several challenges exist due to the unique characteristics of spatiotemporal data. First, spatiotemporal data have various spatial, temporal, and spectral resolutions and diverse sources of noise and errors (e.g., sensor noise, obstacles, and atmospheric effects in remote sensing signals (Licata and Mehta, 2022), GPS errors). Second, spatial registration error and uncertainty may exist when co-registering different layers of geospatial data into the same spatial reference system (He et al., 2022). Third, spatiotemporal data are heterogeneous, i.e., the data distribution often varies across different regions or time periods (Jiang et al., 2019). Thus, a deep learning model trained in one region (or time) may not generalize well to another region (or time). This issue is particularly important when spatial observation samples are sparsely distributed, causing uncertainty when inferring the observations at other locations in continuous space (Hengl et al., 2017; Deng et al., 2023, 2024).
@@
Transportation: Deep learning for transportation data (e.g., ground sensors and video cameras on the road) provides unique opportunities to monitor traffic conditions, analyze traffic patterns, and improve decision-making. For instance, temporal graph neural networks are used to predict traffic flows(e.g., congestion or accidents) (Zhao et al., 2019; Mo and Fu, 2022; Liu et al., 2024b). Deep learning plays a critical role in autonomous driving (e.g., using lidar sensors and optical cameras to detect road lanes and other vehicles or pedestrians). Uncertainty quantification for AI in transportation is challenging due to the temporal dynamics, complex road environment, and the existence of noise and uncertainty (e.g., omission, sparse sensor coverage, errors, or inherent bias). For example, highly crowded events can disrupt normal traffic flows on road networks. Existing works on trajectory prediction uncertainty consider the data uncertainty due to the sparse or insufficient training data (Zhou et al., 2022), and erroneous or missing measurements due to signal loss (Markos et al., 2021). Other works consider complex environmental factors such as extreme weather (Pang et al., 2021; Zhu et al., 2022; Lempert et al., 2022; Wang et al., 2019a). Uncertainty in short-term traffic status forecasting (such as volume, travel speeds, and occupancy) is related to the stochastic environment and model training (Wang et al., 2014), while uncertainty in long-term traffic modeling comes from the exogenous factors of traffic flow (e.g., rainstorms and snowstorms) (Li et al., 2022).